The weather has taken a turn for the worse at LCO. It is currently 1.8C and blowing at 32.7 mph. And getting colder.
Speaking of cold, we installed the new liquid cooling system in the W-Unit today.
Jared and Laird installing the brass cold-plates on our CCDs.
Jason getting ready to install the shutter. Jason first assembled the shutter heatsink and mounting system nearly 2 years ago.
Jared checking the shutter mounting. It is a tight fit, and we don't want any metal-metal contact because the shutter vibrates. A lot.
The team at work on the NAS, under appropriate supervision.
The finished product.
The test gauge for the CCD cooling system.
The other big accomplishment in world of the NAS today was Laird installing the Telecentric Lens. This almost flat lens lets us work off-axis, away from the guide star.
The new telecentric lens.
View past the cube, through the new telecentric lens. That's L1 there.
After installing new glass, we checked the system. Everything looks great - the wavefront is flat and our i' PSF is 3.2 pixels. Click for more pixels.
We also got some time in the dome, and did a little work on the Clay telescope. The vane-end trays are installed, and the leak preventing solenoid valve was plumbed and tested.
Jason and Juan working on the solenoid.
Installing the vane end trays.
This hawk was hanging out over the observatory at lunch today.
A hawk over Magellan. Click for full wingspan.
Tonight’s pretty pictures are from Alan Uomoto, who spent some time messing with his camera’s settings today.
The beginning of the weather. Note the top of the fog in the valley below. Click for bigger.
A panoramic view of the hotel and Las Campanas itself. Click for more awesome.
Days without a motherboard failure: 2